Stock in CDB Industries has a beta of .93. The market risk premium is 7.3 percent, and T-bills are currently yielding 4.3 percent. CDB’s most recent dividend was $2.20 per share, and dividends are expected to grow at a 5.3 percent annual rate indefinitely.

Required:

If the stock sells for $44 per share, what is your best estimate of CDB’s cost of equity? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).)

  Cost of equity %
Answer #1

Cost of equity as per growth model=(D1/Current price)+Growth rate

=[(2.2*1.053)/44]+0.053

=10.565%

Cost of equity as per CAPM=risk free rate+Beta*market risk premium

=4.3+(0.93*7.3)=11.089%

Best estimate of Cost of equity=(10.565+11.089)/2

=10.83%(Approx).